The U.S. Senate consists of 100 members, 2 from each state. A committee of 9 senators is formed. What is the probability that it contains at least one senator from your state?
Answer by stanbon(75887)

The U.S. Senate consists of 100 members, 2 from each state. A committee of 9 senators is formed. What is the probability that it contains at least one senator from your state?
----
P(at least one from my state) = 1 - P(none from my state)
---
= 1 - (98C2)/(100C2)
---
= 1 - 0.96 = 0.04
--------------------
Cheers,
Stan H.
------------------